- In the 1960s, a population of squirrels (Sciurus carolensis) was being studied in Alabama and the coat color of the squirrels was found to range from the more common gray color (dominant) to the less common red color (recessive). When scientists sampled an area, they found 536 gray squirrels and 64 red squirrels. Assuming the population is at Hardy Weinberg’s equilibrium,

In the year 2015, another group of scientists studied the squirrel coat color and found that the population of squirrels had increased. They now found that of 850 squirrels, 115 had red coats.

  1. Has the frequency of the gray and red squirrels changed? If not, why and if they have, why so?
  1. Can you predict whether the frequency of the alleles will change over the next 50 years?
Answer #1

We have to use our chi square test to determine that whether our frequency is changed or not.

Genotype frequency in 1960;

Gray : 536/600= 0.893

Red : 64/600=0.107

Genotype frequency in 2015:

Gray: 850/1000=0.85

Red 115/100=0.135

In terms of 1000 population size.

Chi square test:

Allelic frequency of dominant allele(p) and recessive allele (q) is: